Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
Direct materials 3 feet $ 5 per foot $ 15
Direct labor ? hours ? per hour ?

During March, the company purchased direct materials at a cost of $44,160, all of which were used in the production of 2,350 units of product. In addition, 4,800 direct labor-hours were worked on the product during the month. The cost of this labor time was $36,000. The following variances have been computed for the month:

Materials quantity variance $ 3,150 U
Labor spending variance $ 3,100

U

Labor efficiency variance $ 700

U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.

Req 1A: Actual cost per square foot?
Require 1B: Price variance Spending variance
Require 2C: 2a. For direct labor, compute the standard direct labor rate per hour. (Round your answer to the nearest whole dollar.) 2b. F
Answer #1

1) Material quantity variance = (standard qty-actual qty)Standard price

-3150 = (2350*3*5-5X)

5X = 38400

X(actual qty) = 38400/5 = 7680

Actual Cost per foot = 44160/7680 = 5.75

b) Material price variance = (5-5.75)*7680 = 5760 U

Material spending variance = (2350*15-44160) = 8910 U

2a) Labor rate variance = 3100-700 = 2400 U

Labor rate variance = (Standard rate-actual rate)actual hour

-2400 = (X*4800-36000)

-4800X = -33600

X(Standard rate) = 7

2b) Labor efficiency variance = (Standard hour-actual Hour)Standard rate

-700 = (X*7-4800*7)

-7X = -32900

X(Standard hour) = 4700 Hour

2c) Standard hour per unit = 4700/2350 = 2 hour per unit